# Reading Between the Lines: Towards Reliable Black-box LLM Fingerprinting via Zeroth-order Gradient Estimation

### v1

@@ Appendix B Implementation Details
Detailed Settings of Ablation Studies. For the ablation study of different query data, except the HumanEval dataset (Chen et al., 2021), we select four other datasets as the source of the base queries, including the generative QA dataset TruthfulQA (Lin et al., 2022), the extractive QA dataset SQuAD (Rajpurkar et al., 2016), the complicated reasoning QA dataset DROP (Dua et al., 2019), and a text completion dataset Wikipedia (Foundation, [n. d.]). For the Wikipedia dataset, similar to HumanEval, we sample a text snippet with 20 words and prepend a completion instruction ($i.e.$, “Continue the following text:”) to construct a completion prompt. For the ablation study of different sentence embedding models, we select the following 5 models: all-mpnet-base-v2 (MPNet), all-MiniLM-L6-V2 (MiniLM) (Reimers and Gurevych, 2019), potion-multilingual-128M (Model2Vec) (Tulkens and van Dongen, 2024), Qwen3-Embedding-0.6B (Qwen3) (Zhang et al., 2025a), and EmbeddingGemma-300m (EmbeddingGemma) (Schechter Vera et al., 2025). They are all classic or state-of-the-art sentence embedding models with different embedding dimensions.
@@
Detailed Settings of the Overhead Evaluation. The overhead evaluation is conducted on four derivative models from the Qwen2.5-7B lineage. For fair comparison, we utilize 1 NVIDIA RTX 4090 48GB GPU to compute the runtime of these LLM fingerprinting methods.
Appendix C Ethic Statement

### v2

@@ Appendix B Implementation Details
Detailed Settings of Ablation Studies. For the ablation study of different query data, except the HumanEval dataset (Chen et al., 2021), we select four other datasets as the source of the base queries, including the generative QA dataset TruthfulQA (Lin et al., 2022), the extractive QA dataset SQuAD (Rajpurkar et al., 2016), the complicated reasoning QA dataset DROP (Dua et al., 2019), and a text completion dataset Wikipedia (Foundation, ). For the Wikipedia dataset, similar to HumanEval, we sample a text snippet with 20 words and prepend a completion instruction ($i.e.$, “Continue the following text:”) to construct a completion prompt. For the ablation study of different sentence embedding models, we select the following 5 models: all-mpnet-base-v2 (MPNet), all-MiniLM-L6-V2 (MiniLM) (Reimers and Gurevych, 2019), potion-multilingual-128M (Model2Vec) (Tulkens and van Dongen, 2024), Qwen3-Embedding-0.6B (Qwen3) (Zhang et al., 2025b), and EmbeddingGemma-300m (EmbeddingGemma) (Schechter Vera et al., 2025). They are all classic or state-of-the-art sentence embedding models with different embedding dimensions.
@@
Detailed Settings of the Overhead Evaluation. The overhead evaluation in Section 5.5 is conducted on four derivative models from the Qwen2.5-7B lineage. For fair comparison, we utilize 1 NVIDIA RTX 4090 48GB GPU to compute the runtime of all the evaluated LLM fingerprinting methods.
Table 5. The fingerprint similarities across different LLMs while auditing closed-source LLMs using ZeroPrint.
Type Deriative Independent Models
Models GPT-4o Gemini-2.5 Claude-4.5 DeepSeek-V3
GPT-4o 0.943 0.529 0.570 0.578
Appendix C Case Study of Closed-source Audit
In this section, we include a case study of auditing closed-source LLMs using ZeroPrint. Specifically, we utilize GPT-4o-mini (GPT-4o) as the source model and the same GPT-4o-mini API as the derivative, and take three other LLMs, Gemini-2.5-Flash-Lite (Gemini-2.5), Claude-Haiku-4.5 (Claude-4.5), and DeepSeek-V3, as independent models. The results in Table 5 show that the fingerprint similarity of fingerprints from the same GPT-4o-mini model is $0.943$, which is close to $1$, while other independent LLMs exhibit low similarity ($\approx 0.50$). These results demonstrate the feasibility and potential of our ZeroPrint to audit closed-source, black-box LLMs.
Appendix D Ethic Statement